# Efficient fault-tolerant code switching via one-way transversal CNOT gates

### v1

@@ IV.1 Flag-style code switching
IV.2 Magic state injection
@@ V Scaling up
V.1 Larger-distance codes
@@
Consider, as an example, a QEC code that corrects errors. Say that there exists a single fault in the encoding circuit that causes two errors. For the individual encoding circuit, this is sufficient to call the circuit FT since these two errors will be correctable by the distance-5 code. Now assume that a second fault may happen during the subsequent transversal CNOT such that the total number of errors after both circuits is 3 and can thus lead to failure of the code. In this stricter sense of “sequential fault tolerance”, the single fault in the encoding circuit should at most propagate to a single error, despite . This could be achieved by using flag circuits [15] or Steane-type state preparation, as mentioned in Sec. II.4.
@@
We numerically verify that transversal code switching in both directions indeed works for with the aforementioned codes and the one-way transversal CNOT gates consisting of 17 (19) physical CNOT gates respectively. To prepare both code switching auxiliary qubits fault-tolerantly, we employ non-FT unitary state preparation followed by Steane-type verification (see Fig. 17). In total, the protocols operate with 164 (217) physical qubits and 2317 (3797) CNOT gates999Our non-FT encoding circuits of the state of the () and the state of the () codes have 36 (45) and 159 (298) CNOT gates respectively. Physical qubits from auxiliary state verification are reset and reused where possible. in the typical case where logical auxiliary qubits are verified at the first trial; only 34 (38) CNOTs are used for the actual code switching steps. We show the scaling behavior for the code in Fig. 6. The code switching protocol is only applied to the logical input state, which we expect to be least protected by the triorthogonal code, i.e. (see App. A). It is numerically challenging to collect enough Monte Carlo samples in the low- regime [82]. We stress that, at a realistically attainable entangling gate error rate of , our stabilizer simulations suggest lower logical failure rates for the distance-5 codes, and respectively, as compared to the distance-3 logical -gate, which achieves for the least protected logical input state101010For the three input logical Pauli states on the code, we find failure rates and respectively. We also point out that our scheme seems competitive with Ref. [92] for , where only non-deterministic FT magic state preparation was considered but not the full -gate implementation.. It is expected that the scheme performs slightly worse for the code than for the code due to the larger CNOT gate and qubit overhead. For larger physical error rates , the acceptance rate of logical auxiliary states vanishes quickly for Steane-type state preparation (see Tab. 3). So, it is imperative to devise FT state preparation circuits for larger-distance codes and methods to synthesize such circuits that are practical for the relatively large numbers of physical qubits required by the appropriate 3D codes [76].
V.2 Two-dimensional hardware layout
@@ VI Conclusions and outlook
In all shown code instances, we made use of standard look-up-table decoders, which are known to grow exponentially in size when increasing the code distance. Exploring the use of efficient decoders and the potential of correlated decoding in the context of logical teleportation could be a promising direction for further improving scalability [46, 113, 114, 115]. Possibilities of re-using the logical auxiliary qubits or at least restore their state after the logical teleportation step with potentially fewer physical operations could be a line of future research.
@@
An interesting avenue of future work would be to shed more light on a systematic categorization of appropriate codes for transversal code switching, potentially without the need of explicitly constructing a new doubled code [117]. This should include codes with more than logical qubit. We speculate that this could be possible for surface-code-like QEC codes, which may even include good qLDPC codes [118], such as hypergraph product codes or lift-connected surface codes [119]. The homomorphic CNOT gate, a generalization of the transversal CNOT gate recently employed to perform generalized lattice surgery with encoded auxiliary qubits, could offer a path forward [120, 121, 122].
@@ Acknowledgements
J.H. gratefully acknowledges support by the European Union’s Horizon Europe research and innovation program under Grant Agreement Number 101114305 (“MILLENION-SGA1” EU Project) and the US Army Research Office through Grant Number W911NF-21-1-0007. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of IARPA, the Army Research Office, or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein.
@@ Appendix A Physical quantum circuits
Due to their large size, we provide the stabilizers and encoding circuit of the distance-5 codes in a more compact list format below in order to ease reproducing our results.
@@
Note that these definitions correspond to the original definition of the 3D color code with -type cells and -type faces, which achieves a higher protection against -errors than against -errors, as opposed to the -example discussed in the main text. In simulations for the -case, we use this convention so that the state as input is analogous to an input state with the triorthogonal code definition where - and -type stabilizers are interchanged, such that more - than -errors can be corrected. This interchange does not affect logical failure rate estimations of the state.
@@ Appendix C Measurement-free code switching
Owing to current experimental limitations of quantum computing hardware, there is a growing interest in running FT quantum circuits without measurements of physical qubits [131, 132, 116].
Note that measurement-free code switching basically amounts to performing a SWAP gate between two different QEC codes and discarding the original qubit. A SWAP gate can be decomposed into three CNOT gates, two of which we can perform in the one-way fashion described in the main text. As of now, we are not aware of a unitary logical CNOT gate between the 2D and 3D color code in the other direction. To obtain this “reversed” CNOT gate, we note that one can concatenate each code with the respective other code, e.g., apply encoding circuits of the Tetrahedral code to the Steane code and apply encoding circuits of the Steane code to the Tetrahedral code, which yields a 105-qubit code for both logical qubits. Then, the CNOT gate can be performed transversally between the two logical qubits and decoding circuits of the respective codes can finally be applied to each logical qubit to re-obtain the original code blocks. We conjecture that this could be possible without any additional FT overhead.

### v2

@@ 4.1 Flag-style code switching
4.2 Magic state injection
@@ 5 Scaling up
5.1 Larger-distance codes
@@
Consider, as an example, a QEC code that corrects errors. Say that there exists a single fault in the encoding circuit that causes two errors. For the individual encoding circuit, this is sufficient to call the circuit FT since these two errors will be correctable by the distance-5 code. Now assume that a second fault may happen during the subsequent transversal CNOT such that the total number of errors after both circuits is 3 and can thus lead to failure of the code. In this stricter sense of “sequential fault tolerance”, the single fault in the encoding circuit should at most propagate to a single error, despite . This could be achieved by using flag circuits [15] or Steane-type state preparation, as mentioned in Sec. 2.4.
@@
We numerically verify that transversal code switching in both directions indeed works for with the aforementioned codes and the one-way transversal CNOT gates consisting of 17 (19) physical CNOT gates respectively. To prepare both code switching auxiliary qubits fault-tolerantly, we employ non-FT unitary state preparation followed by Steane-type verification (see Fig. 17). In total, the protocols operate with 164 (217) physical qubits and 2317 (3797) CNOT gates999Our non-FT encoding circuits of the state of the () and the state of the () codes have 36 (45) and 159 (298) CNOT gates respectively. Physical qubits from auxiliary state verification are reset and reused where possible. in the typical case where logical auxiliary qubits are verified at the first trial; only 34 (38) CNOTs are used for the actual code switching steps.
We show the scaling behavior for the code in Fig. 6. The code switching protocol is only applied to the logical input state, which we expect to be least protected by the triorthogonal code, i.e. (see App. A). It is numerically challenging to collect enough Monte Carlo samples in the low- regime [82]. We stress that, at a realistically attainable entangling gate error rate of , our stabilizer simulations suggest lower logical failure rates for the distance-5 codes, and respectively, as compared to the distance-3 logical -gate, which achieves for the least protected logical input state101010For the three input logical Pauli states on the code, we find failure rates and respectively. We also point out that our scheme seems competitive with Ref. [92] for , where only non-deterministic FT magic state preparation was considered but not the full -gate implementation.. It is expected that the scheme performs slightly worse for the code than for the code due to the larger CNOT gate and qubit overhead. For larger physical error rates , the acceptance rate of logical auxiliary states vanishes quickly for Steane-type state preparation (see Tab. 3). So, it is imperative to devise FT state preparation circuits for larger-distance codes and methods to synthesize such circuits that are practical for the relatively large numbers of physical qubits required by the appropriate 3D codes [76].
5.2 Two-dimensional hardware layout
@@ 6 Conclusions and outlook
In all shown code instances, we made use of standard look-up-table decoders, which are known to grow exponentially in size when increasing the code distance. Exploring the use of efficient decoders and the potential of correlated decoding in the context of logical teleportation could be a promising direction for further improving scalability by, for instance, simplifying the logical auxiliary state verification [46, 113, 114, 115]. Additionally, further attention is needed to optimally leverage the potential of using the logical measurements of the teleportation subroutines for intermediate error correction. Possibilities of re-using the logical auxiliary qubits or at least restore their state after the logical teleportation step with potentially fewer physical operations could be another line of future research.
@@
An interesting avenue of future work would be to shed more light on a systematic categorization of appropriate codes for transversal code switching, potentially without the need of explicitly constructing a new doubled code [117]. This should include codes with more than logical qubit. We speculate that this could be possible for surface-code-like QEC codes, which may even include high-rate qLDPC codes [118], such as hypergraph product codes or lift-connected surface codes [119]. The homomorphic CNOT gate, a generalization of the transversal CNOT gate recently employed to perform generalized lattice surgery with encoded auxiliary qubits, could offer a path forward [120, 121, 122].
@@
Author contributions
SH conceptualized the project, analyzed the scheme, designed the circuits and performed all numerical simulations. SH wrote the manuscript with input and feedback from JH.
@@ Acknowledgements
JH gratefully acknowledges support by the European Union’s Horizon Europe research and innovation program under Grant Agreement Number 101114305 (“MILLENION-SGA1” EU Project) and the US Army Research Office through Grant Number W911NF-21-1-0007. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of IARPA, the Army Research Office, or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein.
@@ Appendix A Physical quantum circuits
Due to their large size, we provide the stabilizers and encoding circuits of the 3D distance-5 codes in a more compact list format below in order to ease reproducing our results.
@@
Note that these definitions correspond to the original definition of the 3D color code with -type cells and -type faces, which achieves a higher protection against -errors than against -errors, as opposed to the -example discussed in the main text. In simulations for the -case, we use this convention so that the state as input is analogous to an input state with the triorthogonal code definition where - and -type stabilizers are interchanged, such that more - than -errors can be corrected. This interchange does not affect logical failure rate estimations of the state. All circuits can be accessed under https://doi.org/10.5281/zenodo.15674834.
@@ Appendix C Measurement-free code switching
Owing to current experimental limitations of quantum computing hardware, there is a growing interest in running FT quantum circuits without measurements of physical qubits [131, 132, 116, 133, 134, 135, 136].
Note that measurement-free code switching basically amounts to performing a SWAP gate between two different QEC codes and discarding the original qubit. A SWAP gate can be decomposed into three CNOT gates, two of which we can perform in the one-way fashion described in the main text. As of now, we are not aware of a unitary logical CNOT gate between the 2D and 3D color code in the other direction. To obtain this “reversed” CNOT gate, we note that one can concatenate each code with the respective other code, e.g., apply encoding circuits of the Tetrahedral code to the Steane code and apply encoding circuits of the Steane code to the Tetrahedral code, which yields a 105-qubit code for both logical qubits. Then, the CNOT gate could be performed transversally between the two logical qubits and decoding circuits of the respective codes might finally be applied to each logical qubit to re-obtain the original code blocks. We conjecture that this could be possible without any additional FT overhead.